Re: problem

Date view Thread view Subject view Author view Attachment view

From: Frank Geurts (geurts_at_rice.edu)
Date: Tue May 27 2003 - 13:08:55 EDT


hmm, i see many time-outs but i can't trace back a single attempt of the
shiftcrew to restart the tofpdaq, which might have helped this problem
dissappear ... i restarted the local daq and talked to SCR about it.

-frank

W.J. Llope wrote:
>
>
> Run 4147009 -7 issues - 02:45 TPC anode trip, inner sector #18,
> channel 3. Stop the run 4147007.
> 02:52 TPC recovered, TOF 100% busy and hang the run, removed TOF from RTC.
> 03:00 Starrted run 4147009 successfully. BBC Signal/Background=
> 57/6=9.5, polarization Blue=25%, Yellow=27%
> - Alexei Lebedev
>
>
> guys - we got booted from run control last night (w/ no attempt
> to contact us apparently). frank could you check the daq logs? thanks
>
>
> --
>
>
> W.J. Llope, Ph.D. Res. Assoc. Professor
> T.W. Bonner Nuclear Laboratory phone: 713-348-4741
> Rice University, MS-315 fax: 713-348-5215
> 6100 S. Main St.
> Houston, TX 77005-1892
> e-mail: llope_at_physics.rice.edu
> WWW home: http://wjllope.rice.edu/default.html
>
> --
>


Date view Thread view Subject view Author view Attachment view

This archive was generated by hypermail 2.1.4 : Thu Jul 24 2003 - 00:39:42 EDT